# Improved Treatment of 1-4 interactions in Force Fields for Molecular Dynamics Simulations

### v1

@@ Abstract
Traditional force fields commonly use a combination of bonded torsional terms and empirically scaled non-bonded interactions to capture 1-4 energies and forces of atoms separated by three bonds in a molecule. While this approach can yield accurate torsional energy barriers, it often leads to inaccurate forces and erroneous geometries, and creates an interdependence between dihedral terms and non-bonded interactions, complicating parameterization and reducing transferability. In this paper, we demonstrate that 1-4 interactions can be accurately modeled using only bonded coupling terms, eliminating the need for arbitrarily scaled non-bonded interactions altogether. Furthermore by leveraging the automated parameterization capabilities of the Q-Force toolkit, we efficiently determine the necessary coupling terms without the need for manual adjustment. Our approach is validated on a range of representative molecular systems, encompassing both flexible and rigid structures, and shows a significant improvement in force field accuracy, obtaining sub-kcal/mol mean absolute error for every molecule tested. These results suggest that a bonded-only model for 1-4 interactions can provide a more accurate alternative to that used in traditional force field models for chemical, biomolecular, and materials systems.
@@ 1 Introduction
Accurately simulating molecular systems is fundamental to computational chemistry, enabling advancements in materials science, drug discovery, and biophysics. At the heart of classical simulations are force fields (FFs), physical models that estimate the potential energy of a non-reactive system based on both bonded terms: bond stretching, angle bending, and dihedral torsions, and non-bonded through space interactions including electrostatics, Pauli repulsion and dispersion, and other many-body effects such as polarization and charge transfer. Interactions between atoms separated by a single bond (1-2) or two bonds (1-3) are effectively captured by bond stretching and angle bending terms, respectively. Similarly, interactions between atoms separated by five or more bonds (1-5 and beyond) are typically represented solely by non-bonded terms.
@@
The more common approach, employed by force fields such as OPLS 1, 2, AMBER 3, 4, CHARMM 5, 6, Open Force Field class of FFs 7, 8, and AMOEBA9, 10, is to use a single scaling factor to modulate 1-4 non-bonded interactions. While this method is computationally simple and easy to implement, it is suboptimal for a diverse set of molecular interactions for several reasons. A single scaling parameter cannot capture the nuanced physical phenomena governing 1-4 interactions across different chemical environments, leading to reduced accuracy in force field predictions. While GROMOS 11, 12 incorporates unique 1-4 parameters in specific cases, it defaults to scaled interactions for most systems. While this tailored approach has the potential to improve accuracy for certain molecules or functional groups, it does not fully address the underlying limitations of the standard Lennard-Jones and Coulomb potentials. These functional forms inherently lack the correct physics to represent interactions at short distances characteristic of 1-4 pairs. Although torsional barriers can be improved by fine-tuning 1-4 parameters, this does not guarantee accurate forces or a well-behaved potential energy surface (PES), which are critical for reliable simulations.
To address the inherent limitations of traditional methods for handling 1-4 interactions, we propose two possible pathways for improvement. The first involves treating 1-4 interactions entirely through non-bonded potentials with physically accurate functional forms. This would necessitate the development or adoption of advanced force fields where all non-bonded interactions, including those at short distances, are damped using physically motivated corrections, such as density overlap approaches 13, 14, 15. Force fields like HIPPO 16 and CMM 17, which incorporate density overlap to account for charge penetration and related effects, are promising candidates for this strategy. Such an approach could significantly reduce the reliance on torsional terms by inherently capturing the correct balance of interactions, including electrostatics, van der Waals forces, polarization, and charge transfer. However, while this approach holds promise, achieving the correct balance among these interaction components remains a considerable challenge. Subtle inaccuracies in modeling electrostatics, van der Waals interactions, or polarization effects could introduce errors that torsional terms may still be needed to compensate for. Additionally, the importance of polarization effects, particularly for 1-2 and 1-3 interactions, cannot be overlooked, as they can contribute significantly to the overall polarization energy at the 1-4 level 18. Therefore, even with an advanced non-bonded potential, some torsional corrections may still be necessary to address residual discrepancies and ensure an accurate representation of the energy landscape.
@@
In this work, we revisit the earlier concepts of coupling terms in order to propose a bonded-only treatment for 1-4 interactions, by implementing the coupling terms and developing the necessary protocols for generating and fitting QM data within Q-Force toolkit.25 Q-Force is a framework that enables a systematic, reproducible, and extensible approach to automated force field parameterization based on QM calculations which we harness for defining bonded coupling terms across 1-2, 1-3, and 1-4 coupling terms.
Figure 1: The flexible and rigid molecular datasets used in this study. Molecules in the (a) flexible dataset contains a single fully rotatable torsion, whereas the (b) rigid set has molecules with multiple torsions that are restricted due to double bonds, conjugation or being in a ring.
To validate our approach, we evaluate its performance across a diverse set of molecular systems encompassed in two datasets: the ”flexible set” featuring molecules with flexible torsions (Figure 1a) and the ”rigid set” comprising of rigid molecules with either ring systems or double bonds (Figure 1b). An important choice made here is that we only selected molecules with no discernible non-bonded interactions beyond 1-4, so that we can study the performance of the 1-4 bonded terms without having to choose a specific non-bonded FF, which could affect the interability of the results. We benchmark our results against traditional force fields, including OPLS and Sage, the polarizable force field AMOEBA, and the semi-empirical method xTB 26. Our findings reveal substantial improvements in the accuracy of forces and energy surfaces, demonstrating that a bonded-only treatment of 1-4 interactions can effectively overcome the limitations of hybrid approaches. Additionally, by decoupling the parameterization of torsional and non-bonded terms, our method simplifies force field development, providing a clear and streamlined protocol that can be readily adopted by next-generation of both two-body and advanced force fields.
@@ 2.2 Torsion terms and their coupling
$\textrm{V}_{\textrm{torsion-bond}}(b,\phi)=\sum_{n=1}^{4}K_{n,b\phi}(b-b_{0})(1+\cos(n\phi)$ (8)
@@
$\textrm{V}_{\textrm{torsion-angle}}(\theta,\phi)=\sum_{n=1}^{4}K_{n,\theta\phi}(\cos(\theta)-\cos(\theta_{0})(1+\cos(n\phi))$ (9)
@@ 2.4 QM reference and other benchmark methods
QM reference data was generated using the $\omega$-B97X-V 32 density functional, one of the most accurate hybrid-GGA functionals available 33, 34, in conjunction with the def2-TZVPD 35, basis set. This combination was selected as a good compromise between accuracy and cost. The parameterization protocols in Q-Force, however, are agnostic to the choice of QM method, allowing for the flexibility to adopt more accurate or computationally efficient methods depending on the specific needs of the user or application. Q-Chem software was used for all QM calculations 36.
To benchmark Q-Force against other FFs, we parameterized three additional FFs using well-established tools. The OPLS force field, a widely used standard FF, was parameterized using LigParGen 37. Sage, a modern and QM-based FF developed within the Open Force Field Initiative, was parameterized using OpenForceField tools 38. For a comparison with an advanced FF, we included AMOEBA, a polarizable FF that incorporates various bonded coupling terms10, parameterized using the Poltype2 toolkit 39. These three FFs have 1-4 non-bonded interactions that are scaled by different constants. The xTB GFN2 semi-empirical model26 was also used as an additional benchmark to compare its performance against the FFs, as it uniquely occupies the space between FFs and QM methods in terms of both cost and accuracy. The OpenMM software platform40 was used for all FF evaluations.
We recognize that comparing Q-Force FF, parameterized against the $\omega$-B97X-V reference, with other FFs optimized against different QM methods or objectives, does not constitute a perfectly fair comparison. For instance, Sage is parameterized against B3LYP-D3BJ with DZVP basis set. In order to estimate possible deviations arising from differences in QM references, we included the B3LYP-D3BJ 41, 42, 43 with TZVP basis set as a benchmark. Any error arising from differences in QM references should have similar magnitude to the deviation between these two QM methods. On the other hand, for AMOEBA, we adjusted Poltype settings so that any QM calculations it performs are done with $\omega$-B97X-V with the def2-TZVPD basis set.
@@ 3.1 Energy and force accuracy along the MD trajectory
Figure 2: Absolute energy errors (a) and the norm of the vectorial force error per atom (b) with both datasets combined for all benchmarks in the study. The bin widths are 0.25 kcal/mol for energies and 1.5 $\AA$ for forces. The occurrences are normalized to be 1 at maximum occurrence for each method.
@@
The polarizable AMOEBA FF does better than the fixed charge FFs across energies and forces for the flexible set, but the results are more mixed for the rigid set. OPLS outperforms AMOEBA in energies while AMOEBA does slightly better with forces in this case. AMOEBA’s performance is particularly worse for ring systems, indicating that the automated parameterization strategy of Poltype2 toolkit 39 could be improved in this regard. The semi-empirical xTB’s performance for both energies and forces lies between the FFs and QM methods as expected, with $\sim$1 kcal/mol MAE in energies across both datasets.
@@ 4 Discussion and Conclusions
In contrast, the bonded-only model developed in this work circumvents these issues entirely for modeling 1-4 interactions. By replacing the hybrid treatment of torsional and non-bonded interactions with explicitly parameterized bonded coupling terms, we achieves exceptional accuracy in reproducing QM reference data across a diverse set of molecules, encompassing both flexible and rigid systems. The success of this approach highlights the importance of explicitly accounting for coupled bonded interactions, as we recently showed for bonds and angles27 and now extended to torsion-bond and torsion-angle, torsion-angle-angle terms, which have been underutilized in FF development in our view.
An additional and important feature of this work is the improved development of Q-Force25 for parameterizing these new bond coupling potentials. Notably, the errors in Q-Force predictions were even smaller than the discrepancies between the two QM methods used as benchmarks ($\omega$-B97X-V/def2-TZVPD and B3LYP-D3BJ/TZVP), suggesting that Q-Force closely approaches the inherent accuracy limits of the QM reference method. Importantly, the Q-Force toolkit is agnostic to the QM reference and the same methodology can be used to create FFs parameterized against coupled cluster, MP2 or any other QM method.25 Finally, the automated parameterization capabilities of Q-Force provide a reproducible and scalable framework for FF development in this regard, adaptable to both basic and advanced FFs.
In summary, this work establishes that the bonded-only model is a viable and seemingly better alternative for modeling 1-4 interactions in molecular simulations, and underscores its potential to be used by next-generation FFs using Q-Force. By decoupling torsional and non-bonded parameterization, this methodology simplifies FF development, reduces interdependencies between terms, and will likely enhance transferability. A natural next step is to evaluate the performance of the bonded-only approach in the greater realm of atom-typed force fields, so that they can conveniently be used in larger scale applications such as materials design and drug discovery applications.
@@
S.S. thanks the Dutch Research Council (NWO) Rubicon grant (019.212EN.004) for fellowship support. T.H-G. acknowledges support from the U.S. National Science Foundation through Grant No. CHE-2313791. Computational resources were provided by the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract DE-AC02-05CH11231.

### v2

@@ Abstract
Traditional force fields commonly use a combination of bonded torsional terms and empirically scaled non-bonded interactions to capture 1-4 energies and forces of atoms separated by three bonds in a molecule. While this approach can yield accurate torsional energy barriers, it often leads to inaccurate forces and erroneous geometries, and creates an interdependence between dihedral terms and non-bonded interactions, complicating parameterization and reducing transferability. In this paper, we demonstrate that 1-4 interactions can be accurately modeled using only bonded coupling terms, eliminating the need for arbitrarily scaled non-bonded interactions altogether. Furthermore by leveraging the automated parameterization capabilities of the Q-Force toolkit, we efficiently determine the necessary coupling terms without the need for manual adjustment. Our approach is first validated on a range of small molecule systems, encompassing both flexible and rigid structures, and shows a significant improvement in force field accuracy, obtaining sub-kcal/mol mean absolute error for every molecule tested. We further extend the bonded-only model for 1-4 interactions to Amber ff14sb, CHARMM36, and OPLS-AA force fields to reproduce ab initio gas and implicit solvent $\phi,\psi$ surfaces of alanine dipeptide.
@@ 1 Introduction
Accurately simulating molecular systems is fundamental to computational chemistry, enabling advancements in materials science, drug discovery, and biophysics. At the heart of classical simulations are force fields (FFs), physical models that estimate the potential energy of a (typically) non-reactive system based on both bonded terms: bond stretching, angle bending, and dihedral torsions, and non-bonded through space interactions including electrostatics, Pauli repulsion and dispersion, and other many-body effects such as polarization and charge transfer. Interactions between atoms separated by a single bond (1-2) or two bonds (1-3) are effectively captured by bond stretching and angle bending terms, respectively. Similarly, interactions between atoms separated by five or more bonds (1-5 and beyond) are typically represented solely by non-bonded terms.
@@
The more common approach, employed by force fields such as OPLS 1, 2, AMBER 3, 4, CHARMM 5, 6, the Open Force Field class of FFs 7, 8, and AMOEBA9, 10, is to use a single scaling factor to modulate 1-4 non-bonded interactions. While this method is computationally simple and easy to implement, it is suboptimal for a diverse set of molecular interactions for several reasons. A single scaling parameter cannot capture the nuanced physical phenomena governing 1-4 interactions across different chemical environments, leading to reduced accuracy in force field predictions. While GROMOS 11, 12 incorporates unique 1-4 parameters in specific cases, it defaults to scaled interactions for most systems. While this tailored approach has the potential to improve accuracy for certain molecules or functional groups, it does not fully address the underlying limitations of the standard Lennard-Jones and Coulomb potentials. These functional forms inherently lack the correct physics to represent interactions at short distances characteristic of 1-4 pairs. Although torsional barriers can be improved by fine-tuning 1-4 parameters, this does not guarantee accurate forces or a well-behaved potential energy surface (PES), which are critical for reliable simulations.
To address the inherent limitations of traditional methods for handling 1-4 interactions, we propose two possible pathways for improvement. The first involves treating 1-4 interactions entirely through non-bonded potentials with physically accurate functional forms. This would necessitate the development or adoption of advanced force fields where all non-bonded interactions, including those at short distances, are damped using physically motivated corrections, such as density overlap approaches 13, 14, 15. Force fields like HIPPO 16 and CMM 17, which incorporate density overlap to account for charge penetration and related effects, are promising candidates for this strategy. Such an approach could significantly reduce the reliance on torsional terms by inherently capturing the correct balance of interactions, including electrostatics, van der Waals forces, polarization, and charge transfer. However, while this approach holds promise, achieving the correct balance among these interaction components remains a considerable challenge. Subtle inaccuracies in modeling electrostatics, van der Waals interactions, or polarization effects could introduce errors that may still need to be compensated for by torsional terms. Additionally, the importance of polarization effects, particularly for 1-2 and 1-3 interactions, cannot be overlooked, as they can contribute significantly to the overall polarization energy at the 1-4 level 18. Therefore, even with an advanced non-bonded potential, some torsional corrections may still be necessary to address residual discrepancies and ensure an accurate representation of the energy landscape.
@@
In this work, we revisit the earlier concepts of coupling terms in order to propose a bonded-only treatment for 1-4 interactions, by implementing the coupling terms and developing the necessary protocols for generating and fitting QM data within Q-Force toolkit.25 Q-Force is a framework that enables a systematic, reproducible, and extensible approach to automated force field parameterization based on QM calculations which we harness for defining bonded coupling terms across 1-2, 1-3, and 1-4 coupling terms. We benchmark our results against traditional force fields, including OPLS and Sage, the polarizable force field AMOEBA, and the semi-empirical method xTB 26. Our findings reveal substantial improvements in the accuracy of forces and energy surfaces, demonstrating that a bonded-only treatment of 1-4 interactions can effectively overcome the limitations of hybrid approaches. Additionally, by decoupling the parameterization of torsional and non-bonded terms, our method simplifies force field development, providing a clear and streamlined protocol that can be readily adopted by next-generation of both two-body and advanced force fields. To demonstrate the promise of this approach, we replace the 1-4 nonbonded interactions of AMBER ff14SB, CHARMM36, and OPLS-AA with our bonded-only model for alanine dipeptide in the gas phase and under implicit solvent, and find excellent agreement with the QM data.
@@ 2.2 Torsion terms and their coupling
$V_{torsion-bond}(b,\phi)=\sum_{n=1}^{4}K_{n,b\phi}(b-b_{0})(1+\cos(n\phi)$ (8)
@@
$V_{torsion-angle}(\theta,\phi)=\sum_{n=1}^{4}K_{n,\theta\phi}(cos{\theta}-cos{\theta_{0}}(1+cos{n\phi})$ (9)
@@ 2.3 Parameterization and validation strategy
Figure 1: Flexible and rigid molecular datasets used in this study. Molecules in the (a) flexible dataset contain a single fully rotatable torsion, whereas (b) the rigid set has molecules with multiple torsions that are restricted due to double bonds, conjugation or being in a ring.
To validate our approach, we evaluate its performance across a diverse set of molecular systems encompassed in two datasets: the ”flexible set” featuring molecules with flexible torsions (Figure 1a) and the ”rigid set” comprising of rigid molecules with either ring systems or double bonds (Figure 1b). An important choice made here is that we only selected molecules with no discernible non-bonded interactions beyond 1-4, so that we can study the performance of the 1-4 bonded terms without having to choose a specific non-bonded FF, which could affect the interability of the results.
@@
2.4 Protocol for 2D Torsion scans
For all Ramachandran plots, the minimum energy was subtracted from each point, and only energies within 7 kcal/mol of the minimum were contoured. All energies were graphed using cubic interpolation in python.
  • TorsionDrive32 using xTB GFN2 was used to generate optimized structures at each $\phi$ and $\psi$ point. All angles were held fixed from $165^{\circ}$ to $180^{\circ}$ in $15^{\circ}$ increments. (See Supplementary Figure S4)
  • The optimized coordinates from the xTB 2D scan were used as starting points for QM,AMBER ff14sb, CHARMM36, and OPLS-AA optimizations while holding the corresponding $\phi$ and $\psi$ angles fixed.
  • Converged QM points were used as starting points for Q-Force optimizations at the same fixed $\phi$ and $\psi$ angles.
  • All 1-5 and beyond interactions from AMBER ff14SB, CHARMM36, and OPLS-AA were subtracted from the QM surface, and Q-Force used to fit the 1-2, 1-3, and 1-4 bonded terms.
  • The OpenMM software platform33 was used to perform energy minimizations for all force fields except for OPLS-AA for which GROMOS software package34 was used since OpenMM does not support OPLS-AA natively.
@@ 2.5 QM reference and other benchmark methods
QM reference data was generated using the $\omega$-B97X-V 35 density functional, one of the most accurate hybrid-GGA functionals available 36, 37, in conjunction with the def2-TZVPD 38, basis set. This combination was selected as a good compromise between accuracy and cost. The parameterization protocols in Q-Force, however, are agnostic to the choice of QM method, allowing for the flexibility to adopt more accurate or computationally efficient methods depending on the specific needs of the user or application. Q-Chem software was used for all QM calculations 39.
To benchmark Q-Force against other FFs, we parameterized three additional FFs using well-established tools. The OPLS force field, a widely used standard FF, was parameterized using LigParGen 40. Sage, a modern and QM-based FF developed within the Open Force Field Initiative, was parameterized using OpenForceField tools 41. For a comparison with an advanced FF, we included AMOEBA, a polarizable FF that incorporates various bonded coupling terms10, parameterized using the Poltype2 toolkit 42. These three FFs have 1-4 non-bonded interactions that are scaled by different constants. The xTB GFN2 semi-empirical model26 was also used as an additional benchmark to compare its performance against the FFs, as it uniquely occupies the space between FFs and QM methods in terms of both cost and accuracy. The OpenMM software platform was used for all FF evaluations.
To generate the solvated Ramachandran surfaces we used Solvation Model Density (SMD) for our QM implicit solvent, and GBneck2 (GBN2) for our MM implicit solvent. SMD was chosen because it is a widely used continuum solvent model that leverages the entire solute electron density to accurately describe bulk electrostatic, and non-electrostatic contributions across a variety of solvents. GBN2 was chosen as it is the most recent version of the Generalized Born implicit solvent models that is available in OpenMM/AMBER, and it was designed with proteins and small organic molecules in mind. Their pairing ensures a state-of-the-art approach to implicit solvation models at the QM and MM level. QM energies were computed with the Q-Chem water setting for SMD solvent, while MM energies used GBN2 with $\epsilon=80$.
We recognize that comparing Q-Force FF, parameterized against the $\omega$-B97X-V reference, with other FFs optimized against different QM methods or objectives, does not constitute a perfectly fair comparison. For instance, Sage is parameterized against B3LYP-D3BJ with DZVP basis set. In order to estimate possible deviations arising from differences in QM references, we included the B3LYP-D3BJ 43, 44, 45 with TZVP basis set as a benchmark. Any error arising from differences in QM references should have similar magnitude to the deviation between these two QM methods. On the other hand, for AMOEBA, we adjusted Poltype settings so that any QM calculations it performs are done with $\omega$-B97X-V with the def2-TZVPD basis set.
@@ 3.1 Energy and force accuracy along the MD trajectory
[…]
@@
The polarizable AMOEBA FF does better than the fixed charge FFs across energies and forces for the flexible set, but the results are more mixed for the rigid set. OPLS outperforms AMOEBA in energies while AMOEBA does slightly better with forces in this case. AMOEBA’s performance is particularly worse for ring systems, indicating that the automated parameterization strategy of Poltype2 toolkit 42 could be improved in this regard. The semi-empirical xTB’s performance for both energies and forces lies between the FFs and QM methods as expected, with $\sim$1 kcal/mol MAE in energies across both datasets.
@@ 3.2 Torsional Profiles
3.3 Replacing 1-4 Non-bonded Interactions in Protein Force Fields
To evaluate the efficacy and validity of our new approach to treating 1-4 interactions for any type of force field, we consider the two dimensional $\phi/\psi$ scans of alanine dipeptide. Alanine dipeptide is considered to be a building block of protein models, and the ability to correctly calculate its Ramachandran surface is key to elucidating if a force field is well suited to model conformational changes in proteins. Furthermore, unlike the molecules from the flexible and rigid datasets, alanine dipeptide contains numerous 1-5 and beyond (1-5+) interactions. Hence we use Q-Force to refit the 1-2, 1-3, and 1-4 to the coupled bonded terms and embed them into the AMBER ff14SB, CHARMM36, and OPLS-AA 1-5+ interactions.
Figure 4A shows the QM relaxed $\phi/\psi$ surface of alanine dipeptide in the gas phase. This is compared to the Q-Force derived FFs with the bonded-only functional form, also in the gas phase, as seen in Figures 4B-D. The newly fitted maps reproduce the QM basin locations
Figure 4: Ramachandran $\phi/\psi$ potential energy surfaces for alanine dipeptide. (A) $\omega$-B97X-V/def2-TZVPD, Q-Force for bonded and with (B) ff14sb 1-5+ nonbonded interactions, (C) CHARMM36 nonbonded 1-5+ interactions, (D) OPLS-AA nonbonded 1-5+ interactions. (E) $\omega$-B97X-V/def2-TZVPD with SMD and (F) Q-Force with ff14sb nonbonded interactions and GBN2 implicit solvent; energies are relative to each surface’s minimum to enable the comparison.
and barrier topography very well across the whole $(\phi,\psi)$ surface. Averaged across the 3 different force fields representing the 1-5+ interactions yields a mean absolute error (MAE) $\approx 0.3$ kcal/mole, root mean squared error (RMSE) $\approx 0.5$ kcal/mole, and absolute maximum error of $|\Delta E|\approx 2.25$ kcal/mol at the highest energy points on the map whose energy scale is $\sim$ 100 kcal/mol (see Supplementary Table S3). These deviations arise from the imperfect 1-5+ non-bonded potential from the fixed charge force fields that is used after the bonded fitting, and indicate that they are unable to capture all electronic effects from the QM reference such as polarization, charge transfer, and penetration due to missing many-body interactions.
As previous theory has shown46, the molecular origin of the $\alpha$-helical and left-handed helix minima in the dipeptide Ramanchandran plot is the stabilization that the reaction field from solvent imparts to the aligned peptide dipoles of these configurations, which is a repulsive interaction in the gas phase. This is born out as seen when comparing the gas phase QM result in Figure 4A with the QM and implicit solvent SMD model in Figure 4E. It also allows us to use Q-Force to create a bonded model but now in the presence of implicit solvent as seen in Figure 4F. In the solvated case the agreement with the QM-SMD surface is less good than the gas phase, with MAE $\approx 1.02$ kcal/mole, RMSE $\approx 1.33$ kcal/mole, and max—ΔE— $\approx 3.69$. The major contributor to the larger deviation is in the choice of solvent model employed, i.e., SMD at the QM level versus GBN2 at the MM level.
@@ 4 Discussion and Conclusions
In contrast, the bonded-only model developed in this work circumvents these issues entirely for modeling 1-4 interactions. By replacing the hybrid treatment of torsional and non-bonded interactions with explicitly parameterized bonded coupling terms, we achieve exceptional accuracy in reproducing QM reference data across a diverse set of molecules, encompassing both flexible and rigid systems. The success of this approach highlights the importance of explicitly accounting for coupled bonded interactions, as we recently showed for bonds and angles27 and now extended to torsion-bond and torsion-angle, torsion-angle-angle terms, which have been underutilized in FF development in our view.
An additional and important feature of this work is the improved development of Q-Force25 for parameterizing these new bond coupling potentials. Notably, the errors in Q-Force predictions were even smaller than the discrepancies between the two QM methods used as benchmarks ($\omega$-B97X-V/def2-TZVPD and B3LYP-D3BJ/TZVP), suggesting that Q-Force closely approaches the inherent accuracy limits of the QM reference method. Importantly, the Q-Force toolkit is agnostic to the QM reference and the same methodology can be used to create FFs parameterized against coupled cluster, MP2 or any other QM method.25 Finally, the automated parameterization capabilities of Q-Force provide a reproducible and scalable framework for FF development in this regard as shown with ff14sb, CHARMM36, OPLS-AA, and is adaptable to advanced FFs by the same protocol as well.
In summary, this work establishes that the bonded-only model is a viable and seemingly better alternative for modeling 1-4 interactions in molecular simulations, and underscores its potential to be used by next-generation FFs using Q-Force. By decoupling torsional and non-bonded parameterization, this methodology simplifies FF development, reduces interdependencies between terms, and will enhance transferability. A natural next step is to evaluate the performance of the bonded-only approach in the greater realm of atom-typed force fields, so that they can conveniently be used in larger scale applications such as materials design and drug discovery applications.
@@
S.S. thanks the Dutch Research Council (NWO) Rubicon grant (019.212EN.004) for fellowship support. A.S.A., Y.W. and T.H-G. acknowledges support from the U.S. National Science Foundation through Grant No. CHE-2313791. Computational resources were provided by Savio at UC Berkeley.